OpenStudy (anonymous):

-3x-y=-9: y-4z=7: x-2y+z=-5 how do you solve this

OpenStudy (adamaero):

pick an easy number that fits all three equations and multiply the other two, then subtract

OpenStudy (adamaero):

-3x - y = 9 y - 4z = 7 x -2y + z = -5

OpenStudy (adamaero):

well first you have to deal with only two of the equations, once they're made into one, you do the same to the 3rd

OpenStudy (adamaero):

( y - 4z = 7 ) 4(x -2y + z = -5) _____________________

OpenStudy (adamaero):

y - 4z = 7 4x -8y + 4z = -20 ------------------

OpenStudy (adamaero):

Sometimes you have to subtract and sometimes it's adding the equations so either x, y, or z cancels out

OpenStudy (adamaero):

_____________________ 4x - 7y = -13

OpenStudy (adamaero):

4x - 7y = -13 -3x- y =-9 ______________________ (multiply one equation by something to make x or y equal to the other eq, then add or subtract the whole eq)

OpenStudy (adamaero):

I hope you get it!